# Tighter Confidence Bounds for Sequential Kernel Regression

### v1

@@ D.1 Useful Lemmas
$\forall t\geq 1,\forall x\in\mathcal{X},\qquad\mathrm{LCB}_{\mathcal{F}_{t}}(x)\leq f^{*}(x)\leq\mathrm{UCB}_{\mathcal{F}_{t}}(x).$
@@
$r_{t}$ $=f^{*}(x_{t}^{*})-f^{*}(x_{t})$
$\leq\mathrm{UCB}_{\mathcal{F}_{t-1}}(x_{t}^{*})-\mathrm{LCB}_{\mathcal{F}_{t-1}}(x_{t})$
@@
We use a version of the well-known Elliptical Potential Lemma (see e.g. Lemma 11.11 in (Cesa-Bianchi & Lugosi, 2006), Lemma 5.3 in (Srinivas et al., 2010) and Lemma 11 in (Abbasi-Yadkori et al., 2011)). The version stated in Lemma D.2 is essentially the same as the version in Lemma 5 in (Whitehouse et al., 2023), so we omit the proof. The only difference is that we start with $\sum_{t=1}^{T}\min\left(1,\frac{1}{\alpha}\rho_{\alpha,t-1}^{2}(x_{t})\right)$ instead of $\sum_{t=1}^{T}\frac{1}{\alpha}\rho_{\alpha,t-1}^{2}(x_{t})$, and we use the inequality $\min(1,x)\leq\frac{3}{2}\ln(1+x)$, for all $x\geq 0$ (see Lemma D.8 in (Flynn et al., 2023)) instead of $x\leq 2\ln(1+x)$ for all $x\in[0,1]$. This change means that Lemma D.2 holds for $\alpha>0$, rather than for $\alpha\geq\max(1,\kappa)$, where $\kappa$ is an upper bound on $\sup_{x,x^{\prime}}|k(x,x^{\prime})|$.
Lemma D.2 (Elliptical Potential Lemma).
For any $T\geq 1$ and any $\alpha>0$,
D.2 Proof of Theorem 6.4
Proof of Theorem 6.4.
@@
$\forall t\geq 1,\qquad r_{t}\leq\frac{2\widetilde{R}_{\sigma^{2}/c,t-1}}{\sqrt{\sigma^{2}/c}}\rho_{\sigma^{2}/c,t-1}(x_{t}).$ (27)
From Assumption 6.2 ($f^{*}(x)\in[-C,C]$), we also have
$r_{t}\leq 2C.$ (28)
The combination of (27) and (28) yields
@@
$r_{t}\leq\min\left(2C,\frac{2\widetilde{R}_{\sigma^{2}/c,t-1}}{\sqrt{\sigma^{2}/c}}\rho_{\sigma^{2}/c,t-1}(x_{t})\right)\leq 2\max(C,\widetilde{R}_{\sigma^{2}/c,t-1})\min\left(1,\frac{1}{\sqrt{\sigma^{2}/c}}\rho_{\sigma^{2}/c,t-1}(x_{t})\right).$
Starting with the Cauchy-Schwarz inequality, we have
$\sum_{t=1}^{T}r_{t}\leq\sqrt{T\sum_{t=1}^{T}r_{t}^{2}}\leq\sqrt{T\sum_{t=1}^{T}4\max(C^{2},\widetilde{R}_{\sigma^{2}/c,t-1}^{2})\min\left(1,\frac{1}{\sigma^{2}/c}\rho_{\sigma^{2}/c,t-1}^{2}(x_{t})\right)}.$
From Lemma C.2, we have $\widetilde{R}_{\sigma^{2}/c,t}\leq\sqrt{2\sigma^{2}\gamma_{t}(\sigma^{2}/c)+\tfrac{\sigma^{2}B^{2}}{c}+2\sigma^{2}\ln(\tfrac{1}{\delta})}=:U_{\sigma^{2}/c,t}$. Since $U_{\sigma^{2}/c,t}^{2}$ is monotonically increasing in $t$, we have
$\sum_{t=1}^{T}r_{t}$ $\leq\sqrt{T\sum_{t=1}^{T}4\max(C^{2},U_{\sigma^{2}/c,t-1}^{2})\min\left(1,\frac{c}{\sigma^{2}}\rho_{\sigma^{2}/c,t-1}^{2}(x_{t})\right)}$
$\leq\sqrt{4T\max(C^{2},U_{\sigma^{2}/c,T-1}^{2})\sum_{t=1}^{T}\min\left(1,\frac{c}{\sigma^{2}}\rho_{\sigma^{2}/c,t-1}^{2}(x_{t})\right)}.$
Finally, using the Elliptical Potential Lemma D.2, we have
@@
$\sum_{t=1}^{T}r_{t}\leq\sqrt{12T\gamma_{T}\left(\frac{\sigma^{2}}{c}\right)\max\left(C^{2},2\sigma^{2}\gamma_{T}\left(\frac{\sigma^{2}}{c}\right)+\frac{\sigma^{2}B^{2}}{c}+2\sigma^{2}\ln\left(\frac{1}{\delta}\right)\right)}.$
@@ E.2 Recursive Updates
For a more efficient implementation of our Kernel CMM-UCB, Kernel DMM-UCB and Kernel AMM-UCB algorithms, the inverse $(\boldsymbol{K}_{t}+\alpha\boldsymbol{I})^{-1}$, the log determinant $\ln(\det(\frac{1}{\alpha}\boldsymbol{K}_{t}+\boldsymbol{I}))$ and the right Cholesky factor $\boldsymbol{L}_{t}$ can be updated using the following recursive formulas. These update rules can be derived using properties of the Schur complement (Zhang, 2006), and have been used before in other kernel bandit algorithms (see e.g. App. F of (Chowdhury & Gopalan, 2017)).
@@ F.3 Loose Upper Bounds on $B$ and $\sigma$
To run each kernel bandit algorithm, we only need to know upper bounds on the sub-Gaussian parameter $\sigma$ and the bound $B$ on the RKHS norm of the reward function $f^{*}$. We now investigate what happens when the known bounds on one or both of these quantities are loose.
We run each algorithm in the same synthetic kernel bandit problem as described in Sec. 7, where the kernel is a Matérn kernel with smoothness $\nu=5/2$ and length-scale $\ell=0.5$, and the actions are 3-dimensional vectors in $[0,1]^{3}$. As before, the noise variable are $\sigma$-sub-Gaussian with $\sigma=0.1$ and $\left\|f^{*}\right\|=10$. We run each algorithm with all four pairs of values of $(\sigma,B)$, where $\sigma\in\{0.1,0.2\}$ and $B\in\{10,20\}$.

### v2

@@ D.3 Generic Regret Bound
$\forall t\geq 1,\forall x\in\mathcal{X},\qquad\mathrm{LCB}_{\mathcal{F}_{t}}(x)\leq f^{\star}(x)\leq\mathrm{UCB}_{\mathcal{F}_{t}}(x).$
@@
$r_{t}$ $=f^{\star}(x_{t}^{\star})-f^{\star}(x_{t})$
$\leq\mathrm{UCB}_{\mathcal{F}_{t-1}}(x_{t}^{\star})-\mathrm{LCB}_{\mathcal{F}_{t-1}}(x_{t})$
@@
We can now prove the generic regret bound in Thm. 6.2.
Proof of Thm. 6.2.
@@
$\forall t\geq 1,\qquad r_{t}\leq\frac{2\widetilde{R}_{\sigma^{2}/c,t-1}}{\sqrt{\sigma^{2}/c}}\rho_{\sigma^{2}/c,t-1}(x_{t}).$
From our boundedness assumptions, we also have $r_{t}\leq 2C$. We define the set $\mathcal{T}_{T}(\alpha)$ as
$\mathcal{T}_{T}(\alpha)=\left\{t\in[T]:\frac{1}{\alpha}\rho_{\alpha,t-1}^{2}(x_{t})\geq 1\right\}.$
Recall that $\tau_{T}(\alpha)=|\mathcal{T}_{T}(\alpha)|$. From Lemma B.3, we have $\frac{1}{\alpha}\rho_{\alpha,t-1}^{2}(x_{t})=\|k(\cdot,x_{t})\|_{(\boldsymbol{V}_{t-1}+\alpha\boldsymbol{I}_{\mathcal{H}})^{-1}}^{2}$, which means $\mathcal{T}_{T}(\alpha)$ is the same as the set defined in Lemma 34, with $b=1$. By decomposing the cumulative regret, and then using Cauchy-Schwarz, we obtain
$\sum_{t=1}^{T}r_{t}$ $=\sum_{t\in\mathcal{T}_{T}}r_{t}+\sum_{t\notin\mathcal{T}_{T}}r_{t}$
$\leq 2C\tau_{T}(\tfrac{\sigma^{2}}{c})+\sqrt{T{\textstyle\sum_{t\notin\mathcal{T}_{T}}}r_{t}^{2}}$
@@
$\leq 2C\tau_{T}(\tfrac{\sigma^{2}}{c})+\sqrt{T{\textstyle\sum_{t\notin\mathcal{T}_{T}}}4\widetilde{R}_{\sigma^{2}/c,t-1}^{2}\frac{c}{\sigma^{2}}\rho_{\sigma^{2}/c,t-1}^{2}(x_{t})}$
@@
$\leq 2C\tau_{T}(\tfrac{\sigma^{2}}{c})+\sqrt{4\max_{s\in\{0,1,\dots,T\}}\widetilde{R}_{\sigma^{2}/c,s}^{2}T\sum_{t=1}^{T}\min\left(1,\frac{c}{\sigma^{2}}\rho_{\sigma^{2}/c,t-1}^{2}(x_{t})\right)}.$
Using the bound on the squared radius and Lemma D.4, we have
$\sum_{t=1}^{T}r_{t}\leq 2C\tau_{T}\left(\frac{\sigma^{2}}{c}\right)+\sigma\sqrt{24T\gamma_{T}\left(\frac{\sigma^{2}}{c}\right)\left(\gamma_{T}\left(\frac{\sigma^{2}}{c}\right)+\frac{B^{2}}{2c}+\ln\left(\frac{1}{\delta}\right)\right)}$ (55)
D.4 Regret Bounds By Eigendecay Rate
Here, we determine regret bounds for kernels satisfying the polynomial or exponential eigendecay conditions. We focus on the dependence of the regret on $T$ (and $c$, since we choose this based on $T$). First, suppose that the kernel $k$ satisfies the polynomial eigenvalue decay condition.
Proposition D.13 (Regret Bound for Polynomial Decay).
Suppose that $k$ satisfies the polynomial eigendecay condition, and let $c\propto T^{-\frac{1}{\beta_{p}+1}}$. For any $T\geq 1$ and any $\delta\in(0,1]$, with probability at least $1-\delta$, the cumulative regret of Kernel CMM-UCB, Kernel DMM-UCB (using a grid $A$ containing $\alpha=\sigma^{2}/c$) and Kernel AMM-UCB (using $\alpha=\sigma^{2}/c$) satisfies
$r_{1:T}=\mathcal{O}\left(T^{\frac{\beta_{p}+3}{2\beta_{p}+2}}\ln\left(T\right)^{1-\frac{1}{\beta_{p}}}\right).$
Proof.
From the bounds in Lemma D.3 and Lemma D.10, we have
@@ D.4 Regret Bounds By Eigendecay Rate
$\gamma_{T}\left(\tfrac{\sigma^{2}}{c}\right)=\mathcal{O}\left(\left(cT\right)^{\frac{1}{\beta_{p}}}\ln\left(cT\right)^{1-\frac{1}{\beta_{p}}}\right),~{}~{}\text{and}~{}~{}\tau_{T}\left(\tfrac{\sigma^{2}}{c}\right)=\mathcal{O}\left(c^{\frac{1}{\beta_{p}-1}}\ln\left(c^{\frac{\beta_{p}}{\beta_{p}-1}}\right)\right).$
If we plug these into the regret bound in Eq. (55), we get
$\sum_{t=1}^{T}r_{t}=\mathcal{O}\left(c^{\frac{1}{\beta_{p}-1}}\ln\left(c^{\frac{\beta_{p}}{\beta_{p}-1}}\right)+T^{\frac{1}{2}+\frac{1}{\beta_{p}}}c^{\frac{1}{\beta_{p}}}\ln(cT)^{1-\frac{1}{\beta_{p}}}+T^{\frac{1}{2}+\frac{1}{2\beta_{p}}}c^{\frac{1}{2\beta_{p}}-\frac{1}{2}}\ln(cT)^{\frac{1}{2}-\frac{1}{2\beta_{p}}}\right).$
If we set $c\propto T^{q}$, for some $q$ to be chosen later, this becomes
$\sum_{t=1}^{T}r_{t}=\mathcal{O}\left(T^{\frac{q}{\beta_{p}-1}}\ln\left(T^{\frac{q\beta_{p}}{\beta_{p}-1}}\right)+T^{\frac{1}{2}+\frac{1}{\beta_{p}}+\frac{q}{\beta_{p}}}\ln\left(T^{1+q}\right)^{1-\frac{1}{\beta_{p}}}+T^{\frac{1}{2}+\frac{1}{2\beta_{p}}+\frac{q}{2\beta_{p}}-\frac{q}{2}}\ln\left(T^{1+q}\right)^{\frac{1}{2}-\frac{1}{2\beta_{p}}}\right).$
If we choose $q=-\frac{1}{\beta_{p}+1}$, then
$\frac{1}{2}+\frac{1}{\beta_{p}}+\frac{q}{\beta_{p}}=\frac{1}{2}+\frac{1}{2\beta_{p}}+\frac{q}{2\beta_{p}}-\frac{q}{2}=\frac{\beta_{p}+3}{2\beta_{p}+2}.$
Therefore, with this choice of $q$, the regret bound is
$\sum_{t=1}^{T}r_{t}=\mathcal{O}\left(T^{\frac{\beta_{p}+3}{2\beta_{p}+2}}\ln\left(T\right)^{1-\frac{1}{\beta_{p}}}\right).$
This matches the regret bound in Thm. 2 in (Whitehouse et al.,, 2023), which we believe to be the best known regret bound for (vanilla) KernelUCB/GP-UCB under polynomial eigendecay. It is known that the Matérn kernel, with smoothness parameter $\nu>1/2$, satisfies the polynomial eigendecay condition with $\beta_{p}=\frac{2\nu+d}{d}$ (Santin and Schaback,, 2016). Therefore, if we choose $c\propto T^{-\frac{d}{2\nu+2d}}$, the the regret bound is
$\sum_{t=1}^{T}r_{t}=\mathcal{O}\left(T^{\frac{\nu+2d}{2\nu+2d}}\ln\left(T\right)^{\frac{2\nu}{2\nu+d}}\right).$
Now, suppose that the kernel $k$ satisfies the exponential eigendecay condition.
Proposition D.14 (Regret Bound for Exponential Eigendecay).
Suppose that $k$ satisfies the exponential eigendecay condition, and let $c\propto 1$. For any $T\geq 1$ and any $\delta\in(0,1]$, with probability at least $1-\delta$, the cumulative regret of Kernel CMM-UCB, Kernel DMM-UCB (using a grid $A$ containing $\alpha=\sigma^{2}/c$) and Kernel AMM-UCB (using $\alpha=\sigma^{2}/c$) satisfies
$r_{1:T}=\mathcal{O}\left(\sqrt{T}\ln\left(T\right)^{1+\frac{1}{\beta_{e}}}\right).$
Proof.
From the bounds in Lemma D.3 and Lemma D.11, we have
$\gamma_{T}\left(\tfrac{\sigma^{2}}{c}\right)=\mathcal{O}\left(\ln\left(cT\right)^{1+\frac{1}{\beta_{e}}}\right),~{}~{}\text{and}~{}~{}\tau_{T}\left(\tfrac{\sigma^{2}}{c}\right)=\mathcal{O}\left(\max\left(\tfrac{1}{c},\ln\left(c\right)^{1+\frac{1}{\beta_{e}}}\right)\right).$
If we plug these into Eq. (55), we get
$\sum_{t=1}^{T}r_{t}=\mathcal{O}\left(\max\left(\tfrac{1}{c},\ln\left(c\right)^{1+\frac{1}{\beta_{e}}}\right)+\sqrt{T}\ln\left(cT\right)^{1+\frac{1}{\beta_{e}}}+\sqrt{\tfrac{T}{c}\ln\left(cT\right)^{1+\frac{1}{\beta_{e}}}}\right).$
If we set $c\propto 1$, this becomes
$\sum_{t=1}^{T}r_{t}=\mathcal{O}\left(\sqrt{T}\ln\left(T\right)^{1+\frac{1}{\beta_{e}}}\right).$
It is known that the RBF kernel satisfies the exponential eigendecay condition with $\beta_{e}=1/d$ (Belkin,, 2018). This means that, when $c\propto 1$, the regret bound is
$\sum_{t=1}^{T}r_{t}=\mathcal{O}\left(\sqrt{T}\ln\left(T\right)^{d+1}\right).$
Note that one could also use the regret bound in (55) to determine how $c$ should depend on $B$. For exponential eigendecay, one can show that the choice $c\propto B^{2}$ leads to a regret bound which is polylogarithmic in $B$ in the dominant $\sqrt{T}$ term. One can also consider the dependence of the regret on $\sigma$. For exponential eigendecay, when $\sigma\propto 1/\sqrt{T}$, the regret bound becomes polylogarithmic in $T$. Therefore, KernelUCB with our confidence bounds can still enjoy $\mathrm{polylog}(T)$ regret even when the problem is only approximately noiseless.
D.5 Regret Bounds for the Noiseless Setting
In Sec. 6.3, we showed that for all $\alpha>0$ and $T\geq 1$,
$r_{1:T}\leq 2C\tau_{T}(\alpha)+\sqrt{12\alpha B^{2}T\gamma_{T}(\alpha)}.$ (56)
First, suppose that the kernel $k$ satisfies the polynomial eigendecay condition. In the noiseless setting, we can prove the following regret bound.
Proposition D.15 (Regret Bound for Polynomial Decay in the Noiseless Setting).
Suppose that $k$ satisfies the polynomial eigendecay condition, and let $\alpha\propto T^{\frac{1-\beta_{p}^{2}}{\beta_{p}^{2}+1}}$. For any $T\geq 1$, the cumulative regret of Kernel CMM-UCB, Kernel DMM-UCB (using a grid $A$ containing this $\alpha$) and Kernel AMM-UCB (using this $\alpha$) satisfies
$r_{1:T}=\mathcal{O}\left(T^{\frac{\beta_{p}+1}{\beta_{p}^{2}+1}}\ln\left(T\right)\right).$
This is always sublinear in $T$, since $\beta_{p}>1$, and has a growth rate slower than $\sqrt{T}$ whenever $\beta_{p}>1+\sqrt{2}$.
Proof.
We’re interested in the case where $\alpha\propto T^{q}$ for some $q<0$. In this regime, the bounds in Lemma D.3 and Lemma D.10 tell us that
$\tau_{T}(\alpha)=\mathcal{O}\left(\alpha^{-\frac{1}{\beta_{p}-1}}\ln\left(\alpha^{-\frac{\beta_{p}}{\beta_{p}-1}}\right)\right),\quad\gamma_{T}(\alpha)=\mathcal{O}\left(\left(\tfrac{T}{\alpha}\right)^{\frac{1}{\beta_{p}}}\ln\left(\tfrac{T}{\alpha}\right)^{1-\frac{1}{\beta_{p}}}\right).$
This means that Eq. (56) becomes
$r_{1:T}=\mathcal{O}\left(\alpha^{-\frac{1}{\beta_{p}-1}}\ln\left(\alpha^{-\frac{\beta_{p}}{\beta_{p}-1}}\right)+T^{\frac{1}{2}+\frac{1}{2\beta_{p}}}\alpha^{\frac{1}{2}-\frac{1}{2\beta_{p}}}\ln\left(\tfrac{T}{\alpha}\right)^{\frac{1}{2}-\frac{1}{2\beta_{p}}}\right).$
If $\alpha\propto T^{q}$, for some $q<0$ to be chosen later, then
$r_{1:T}=\mathcal{O}\left(T^{-\frac{q}{\beta_{p}-1}}\ln\left(T^{-\frac{\beta_{p}q}{\beta_{p}-1}}\right)+T^{\frac{1}{2}+\frac{1}{2\beta_{p}}+\frac{q}{2}-\frac{q}{2\beta_{p}}}\ln\left(T^{1-q}\right)^{\frac{1}{2}-\frac{1}{2\beta_{p}}}\right).$
We want to choose $q$ such that the maximum of the powers of $T$ is minimised. We notice that, since $\beta_{p}>1$, $-\frac{q}{\beta_{p}-1}$ is monotonically decreasing in $q$ and $\frac{1}{2}+\frac{1}{2\beta_{p}}+\frac{q}{2}-\frac{q}{2\beta_{p}}$ is monotonically increasing in $q$. If we choose $q=\frac{1-\beta_{p}^{2}}{\beta_{p}^{2}+1}$, then
$-\frac{q}{\beta_{p}-1}=\frac{1}{2}+\frac{1}{2\beta_{p}}+\frac{q}{2}-\frac{q}{2\beta_{p}}=\frac{\beta_{p}+1}{\beta_{p}^{2}+1}.$
Therefore, with this choice of $q$, we have
$r_{1:T}=\mathcal{O}\left(T^{\frac{\beta_{p}+1}{\beta_{p}^{2}+1}}\ln\left(T\right)\right).$
Consider the the Matérn kernel, where $\beta_{p}=\frac{2\nu+d}{d}$. If we choose $\alpha\propto T^{-\frac{2\nu^{2}+2\nu d}{2\nu^{2}+2\nu d+d^{2}}}$, then the regret bound for the noiseless case is
$r_{1:T}=\mathcal{O}\left(T^{\frac{\nu d+d^{2}}{2\nu^{2}+2\nu d+d^{2}}}\ln\left(T\right)\right).$
Whenever $\nu>d/\sqrt{2}$, the growth rate is slower than $\sqrt{T}$. Now, suppose that the kernel satisfies the exponential eigendecay condition. In the noiseless setting, we can prove the following regret bound.
Proposition D.16 (Regret Bound for Exponential Decay in the Noiseless Setting).
Suppose that $k$ satisfies the exponential eigendecay condition, and let $\alpha\propto 1$. For any $T\geq 1$, the cumulative regret of Kernel CMM-UCB, Kernel DMM-UCB (using a grid $A$ containing this $\alpha$) and Kernel AMM-UCB (using this $\alpha$) satisfies
$r_{1:T}=\mathcal{O}\left(\ln\left(T\right)^{1+\frac{1}{\beta_{e}}}\right).$
Proof.
We’re interested in the case where $\alpha\propto T^{q}$ for some $-1\leq q\leq 0$. In this regime, the bounds in Lemma D.3 and Lemma D.11 tell us that
$\tau_{T}(\alpha)=\mathcal{O}\left(\ln\left(\tfrac{1}{\alpha}\right)^{1+\frac{1}{\beta_{e}}}\right),\quad\gamma_{T}(\alpha)=\mathcal{O}\left(\ln\left(\tfrac{T}{\alpha}\right)^{1+\frac{1}{\beta_{e}}}\right).$
This means that Eq. (56) becomes
$r_{1:T}=\mathcal{O}\left(\ln\left(\tfrac{1}{\alpha}\right)^{1+\frac{1}{\beta_{e}}}+T^{\frac{1}{2}}\alpha^{\frac{1}{2}}\ln\left(\tfrac{T}{\alpha}\right)^{\frac{1}{2}+\frac{1}{2\beta_{e}}}\right).$
If we choose $\alpha=1/T$, then
$r_{1:T}=\mathcal{O}\left(\ln\left(T\right)^{1+\frac{1}{\beta_{e}}}\right).$
For the RBF kernel, where $\beta_{e}=1/d$, we get
$r_{1:T}=\mathcal{O}\left(\ln\left(T\right)^{d+1}\right).$
@@ E.2 Recursive Updates
For a more efficient implementation of our Kernel CMM-UCB, Kernel DMM-UCB and Kernel AMM-UCB algorithms, the inverse $(\boldsymbol{K}_{t}+\alpha\boldsymbol{I})^{-1}$, the log determinant $\ln(\det(\frac{1}{\alpha}\boldsymbol{K}_{t}+\boldsymbol{I}))$ and the right Cholesky factor $\boldsymbol{L}_{t}$ can be updated using the following recursive formulas. These update rules can be derived using properties of the Schur complement (Zhang,, 2006), and have been used before in other kernel bandit algorithms (see e.g. App. F of (Chowdhury and Gopalan,, 2017)).
@@ F.3 Loose Upper Bounds on $B$ and $\sigma$
To run each kernel bandit algorithm, we only need to know upper bounds on the sub-Gaussian parameter $\sigma$ and the bound $B$ on the RKHS norm of the reward function $f^{\star}$. We now investigate what happens when the known bounds on one or both of these quantities are loose.
We run each algorithm in the same synthetic kernel bandit problem as described in Sec. 7, where the kernel is a Matérn kernel with smoothness $\nu=5/2$ and length-scale $\ell=0.5$, and the actions are 3-dimensional vectors in $[0,1]^{3}$. As before, the noise variable are $\sigma$-sub-Gaussian with $\sigma=0.1$ and $\left\|f^{\star}\right\|=10$. We run each algorithm with all four pairs of values of $(\sigma,B)$, where $\sigma\in\{0.1,0.2\}$ and $B\in\{10,20\}$.